OpenStudy (anonymous):

Calculate the concentration of [H3O+] for each of the pH pH=12 pH=1 pH=7 Dont really understand how to calculate these? Help?

OpenStudy (jfraser):

pH and concentration are related by a pair of equations: \[ph = -\log[H_3O^{+1}]\]and\[[H_3O^{+1}] = 10^{-pH}\]you'll need a calculator for more complicated problems, but at the moment all you need to do is substitute the pH in the second equation with the given pH, and simplify\[[H_3O^{+1}] = 10^{-pH} = 10^{-7} = 1*10^{-7}M = 0.0000007M\]